OpenStudy (anonymous):

4x*2+4x+1=0 using the ax*2+bx+c=0 of the quadratic equation above factor the left hand side of the equation into two linear factors

myininaya (myininaya):

4x^2+2x+2x+1=0 2x(2x+1)+1(2x+1)=0 (2x+1)(2x+1)=0 (2x+1)^2=0